0.96g of H ydrogen iodide was heated to 450 degree centigrade till the equilibrium was reached.It was then quickly cooled and the amount of iodine formed required 15.7ml of N/10 Na 2 S 2 O 3 . Calculate the degree of dissociation of HI at 450 degree centigrade

0.96g of Hydrogen iodide was heated to 450 degree centigrade till the equilibrium was reached.It was then quickly cooled and the amount of iodine formed required 15.7ml of N/10 Na2S2O3. Calculate the degree of dissociation of HI at 450 degree centigrade

Grade:12

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

I will give u a hint regarding this:

The degree of dissociation of PCl5 at a certain temperature and under atmospheric pressure is 0.2. Calculate the pressure at which it will be half dissociated at the same temperature.

PCl5 dissociates as:
PCl5 ? PCl3 Cl2
If α is the degree of dissociation at certain temperature under atmospheric pressure, then
Initial concentration:
PCl5 = 1
PCl3 = 0
Cl2 = 0
At Equilibrium:
PCl5 = 1 – α
PCl3 = α
Cl2 = α
Total number of moles at equilibrium = 1 – α + α + α = 1 + α
Partial pressures of PCl5, PCl3 and Cl2 will be:
p (PCl3) = α p / 1 + α
p (Cl2) = α p / 1 + α
p (PCl5) = (1 – α) p / 1 + α
Kp = p (PCl3) X p (Cl2) / p (PCl5)
Kp = [(α p / 1 + α) X (α p / 1 + α)]/[ (1 – α) p / (1 + α)]
= α2 p / (1 – α) 2
Substituting p = 1 atm and α = 0.2
Kp = (0.2) 2 X 1 / (1 – (0.2)) 2
= 0.041
When α = 1/2 = 0.5, let pressure is p
Kp = α2 p / (1 – α) 2
0.041 = (0.5) sup>2 p / (1 – (0.5)) 2
p = (0.041) [1 – (0.5) 2] / (0.5) 2
p = 0.125 atm

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi